OpenStudy (anonymous):

Explain how raising a quotient to a power is similar to raising a product to a power

OpenStudy (jim766):

product to a power (3x)^2 both the 3 and the x get squared 3^2x^2 = 9x^2 quotient to a power (3/x)^2 both the numerator and the denominator get raised to the 2nd power 3^2/x^2 = 9/x^2 make sense?